Knight, Inc., has issued a three-year bond that pays a coupon of 6.84 percent. Coupon payments are made semiannually. Given the market rate of interest of 4.97 percent, what is the market value of the bond? (Round answer to 2 decimal places, e.g. 15.25.)
Answer #1

Price of Bond = C *[( 1 - ( 1 +R)^-N)] / R + FV / ( 1 +R)^N

Where, C = Coupon payment

R = Rate of Interest of market per period

N = Number of period

FV = face value

Coupon payment = Face value * Coupon rate * ( 1 / Number of compounding in a year)

Face Value = 1000 [ Assumed]

Coupon Payment = 1000 * 6.84% * ( 1/2)

= 34.2

Market interest rate per period = 4.97%/2

= 2.485%

Number of Periods = 3*2

= 6

So, Market value of bond = 34.2 * [( 1 - ( 1 +2.485%)^-6) / 2.485%] + 1000 / ( 1 + 2.485%)^6

= 34.2 * [( 1 - 0.863054392) / 0.02485] + 1000 / 1.15867552329

= 34.2 * 5.51088965795 + 863.054392623

= 1051.5268189

So, Bond value = 1051.5268189